Find the equation of the circle concentric with the circle
x*2+y*2- 6x+ 2y- 11 = 0 and passing through the centre of the circle
x*2+y*2= 16.

Answer by Boreal(15235)

rewrite the first as
x^2-6x+y^2+2y=11
complete the square
x^2-6x+9+y^2+2y+1=21
(x-3)^2+(y+1)^2=21
first circle has center at (3, -1) and radius sqrt (21)
The circle concentric with this has center (3, -1) and passes through the origin, since the center of x^2+y^2=16 is a circle with center (0, 0) and radius 4
The radius of the second circle is the distance from the center to the origin, which is sqrt (10).
Its equation is (x-3)^2+(y+1)^2=10 ANSWER